A Bibliography of Papers in Lecture Notes in Computer Science (2011)

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: http://www.math.utah.edu/~beebe/

14 October 2017 Version 1.03

Title word cross-reference

t [18].

-Cheater [18].

5th [57].

Account [51]. ACS [50]. Adaptive [14]. Adversaries [23]. AES [7]. Algorithms [22, 3]. Almost [18]. Amplification [30]. Applications [14, 2]. April [57]. Arithmetic [2]. Assumption [26]. Atomicity [40]. Attack [21]. Attacks [19, 9]. Attribute [32, 31]. Attribute-Based [32, 31]. Authentication [16, 4]. Autonomic [45]. Avatar [38].

Back [35]. Barrier [20]. Based [32, 31, 26]. Belgian [51]. Beyond [20]. Biographical [37]. Birthday [20]. Blocks [54]. Bounded [27]. Brian [48, 49, 36, 37]. Building [38].

CA [56]. Careful [28]. Case [5]. Channel [17, 9]. Cheater [18]. Chemical [45]. Circuit [29]. Circuit-Size [29]. Closet [46]. Cloud [40]. Codex [2]. Commitments [26]. communication [57]. Commuting [15]. Compact [7]. Completeness [30]. Composable [26]. Composition [28, 27]. Computation [25, 12, 24, 23, 57]. Computer [46, 49]. Computing [6, 45, 48, 42]. Concurrent [27]. conference [56, 57]. Congestion [44]. Construction [34]. Criticality [47]. Cryptanalysis [19]. cryptographer [56]. cryptography [57]. cryptology [56]. Cryptosystems [33]. CT [56]. CT-RSA [56]. Curves [6].

Data [44]. DDH [26]. Decentralizing [32]. Deniable [34]. Dependability [52]. Dependent [30]. Detection [34]. Development [55]. Devices [9]. Domain [20]. DSS [41].

Early [46]. Efficient [24, 4, 26, 29]. Electronic [51, 42]. Encryption [12, 34, 15, 10, 32, 31, 29]. Evaluate [44]. Explicit [6]. Extension [20]. Extractable [33].

Faster [6]. Fault [45, 53]. Faults [47]. February [56]. Formal [9]. Formulas [6]. Fragmentation [49]. Framework [52, 28]. Francisco [56]. free [14]. Front [1]. Fully [8, 10]. Fully-Homomorphic [10]. Functions [11].

Game [25]. Generic [30, 22]. Gentry [10]. Groups [14].

Hard [22, 4]. Hash [33]. HIBE [31]. Highly [26]. Highly-Efficient [26]. History [48]. Homomorphic [11, 10, 12]. Hulls [19].

IBM [50]. IBM-ACS [50]. Ideal [5]. Identifiable [18]. Implementation [7]. Implementing [10]. Improved [22]. Independent [29]. Indifferentiability [28]. Information [16]. Interactive [34, 24]. Intrusion [52]. Invention [39]. Issues [9].

KDM [29]. Key [30, 16, 29]. Key-Dependent [30]. Keys [17]. Knapsacks [22].

Lattice [3]. Lattices [5]. Leakage [8]. Leakage-Resilient [8]. Leaking [16].

Learned [50, 36]. Learning [4]. Leeds [57]. Lessons [50]. Limitations [28]. Limits [7]. Linear [19]. Local [40].

Machine [51]. MACs [20]. Making [5]. Malicious [23]. Mathematical [51]. Matter [35, 1]. Message [30]. MILS [41]. Mixed [47]. Model [27]. Multiparty [12]. Museum [46].

Nanoscale [9]. Negligible [34]. Noise [17]. Non [24]. Non-interactive [24]. Note [37]. NTRU [5].

Optimum [18]. Oracles [13]. Ordinary [6]. Output [23].

Pairings [6]. papers [57]. Pioneering [49]. Polynomial [11]. Power [9]. Practice [39, 3]. Pre [42]. Pre-electronic [42]. PRESENT [19]. Probability [34]. Problems [4, 5]. proceedings [56]. Process [55]. Professor [48]. Programming [39]. Project [50]. Proofs [13, 33]. Pseudo [14]. Pseudo-free [14]. Public [29]. PUFFIN [19]. Pushing [7].

Quantum [27, 57].

Randell [48, 49, 37]. Random [13]. RC4 [21]. Real [44]. Real-Time [44]. Recovery [54]. Reduction [3]. Referee [38]. Reminiscences [50]. Resilience [52]. Resilient [8]. revised [57]. Revocation [33]. Road [44]. RSA [56].

San [56]. Saturation [19]. Scheme [10]. Schemes [18, 13]. Secret [17, 18]. Secure [25, 16, 24, 5, 23]. Security [30, 52, 29]. selected [57]. Semi [12]. Semi-homomorphic [12]. Setting [45]. Sharing [18]. Side [9]. Side-Channel [9]. Signature [13]. Signatures [11, 8, 15]. Size [29]. Statistical [19, 21]. Storage [49, 27].

Structuring [55]. Study [9]. Supercomputer [50]. Systems [45, 47].

Theoretic [25]. Theory [39, 2, 3, 57]. Threats [52]. Threshold [7, 33]. Tight [13]. Time [44]. Timing [47]. Tolerance [53]. Tolerant [45]. Topics [56]. TQC [57]. track [56]. Traffic [44]. Transactions [40]. Two [23]. Two-Output [23].

UK [57]. Unbounded [31]. Universally [26]. Universally-Composable [26]. USA [56]. Using [44].

Variability [9]. Verifiable [15]. Very [7]. via [33]. View [25]. Virtual [53].

Wanderings [43]. Weak [16]. Whetstone [43]. without [16, 13]. Work [49]. Worst [5]. Worst-Case [5]. Writing [55].

Years [46].

References

Anonymous:2011:FM

[1] Anonymous. Front matter. Lecture Notes in Computer Science, 6632:??, 2011. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/content/pdf/bfm:
978-3-642-20465-4/1.

Cramer:2011:ACT

[2] Ronald Cramer. The arithmetic codex: Theory and applications. Lecture Notes in Computer Science, 6632:1, 2011. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/content/pdf/10.
1007/978-3-642-20465-4_1.

Nguyen:2011:LRA

[3] Phong Q. Nguyen. Lattice reduction algorithms: Theory and practice. Lecture Notes in Computer Science, 6632: 2–6, 2011. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer.com/content/
pdf/10.1007/978-3-642-20465-4_2.

Kiltz:2011:EAH

[4] Eike Kiltz, Krzysztof Pietrzak, David Cash, Abhishek Jain, and Daniele Venturi. Efficient authentication from hard learning problems. Lecture Notes in Computer Science, 6632:7–26, 2011. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/content/pdf/10.
1007/978-3-642-20465-4_3.

Stehle:2011:MNS

[5] Damien Stehlé and Ron Steinfeld. Making NTRU as secure as worst-case problems over ideal lattices. Lecture Notes in Computer Science, 6632:27–47, 2011. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer.com/content/
pdf/10.1007/978-3-642-20465-4_4.

Aranha:2011:FEF

[6] Diego F. Aranha, Koray Karabina, Patrick Longa, Catherine H. Gebotys, and Julio López. Faster explicit formulas for computing pairings over ordinary curves. Lecture Notes in Computer Science, 6632:48–68, 2011. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/content/pdf/10.
1007/978-3-642-20465-4_5.

Moradi:2011:PLV

[7] Amir Moradi, Axel Poschmann, San Ling, Christof Paar, and Huaxiong Wang. Pushing the limits: A very compact and a threshold implementation of AES. Lecture Notes in Computer Science, 6632:69– 88, 2011. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer.com/content/
pdf/10.1007/978-3-642-20465-4_6.

Boyle:2011:FLR

[8] Elette Boyle, Gil Segev, and Daniel Wichs. Fully leakage-resilient signatures. Lecture Notes in Computer Science, 6632:89–108, 2011. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.
com/content/pdf/10.1007/978-3-642-
20465-4_7.

Renauld:2011:FSP

[9] Mathieu Renauld, François-Xavier Standaert, Nicolas Veyrat-Charvillon, and Dina Kamel. A formal study of power variability issues and side-channel attacks for nanoscale devices. Lecture Notes in Computer Science, 6632:109– 128, 2011. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer.com/content/
pdf/10.1007/978-3-642-20465-4_8.

Gentry:2011:IGF

[10] Craig Gentry and Shai Halevi. Implementing Gentry’s fully-homomorphic encryption scheme. Lecture Notes in Computer Science, 6632:129–148, 2011. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/content/pdf/10.
1007/978-3-642-20465-4_9.

Boneh:2011:HSP

[11] Dan Boneh and David Mandell Freeman. Homomorphic signatures for polynomial functions. Lecture Notes in Computer Science, 6632:149–168, 2011. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

                                        /link.springer.com/content/pdf/10.
                                        1007/978-3-642-20465-4_10.

Bendlin:2011:SHE

[12] Rikke Bendlin, Ivan Damgaård, Claudio Orlandi, and Sarah Zakarias. Semihomomorphic encryption and multiparty computation. Lecture Notes in Computer Science, 6632:169–188, 2011. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

                                        /link.springer.com/content/pdf/10.
                                        1007/978-3-642-20465-4_11.

Schage:2011:TPS

[13] Sven Schäge. Tight proofs for signature schemes without random oracles. Lecture Notes in Computer Science, 6632:189– 206, 2011. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

                                        http://link.springer.com/content/
                                        pdf/10.1007/978-3-642-20465-4_12.

Catalano:2011:APF

[14] Dario Catalano, Dario Fiore, and Bogdan Warinschi. Adaptive pseudo-free groups and applications. Lecture Notes in Computer Science, 6632:207–223, 2011. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

                                        /link.springer.com/content/pdf/10.
                                        1007/978-3-642-20465-4_13.

Fuchsbauer:2011:CSV

[15] Georg Fuchsbauer. Commuting signatures and verifiable encryption. Lecture Notes in Computer Science, 6632:224– 245, 2011. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer.com/content/
pdf/10.1007/978-3-642-20465-4_14.

Bouman:2011:SAW

[16] Niek J. Bouman and Serge Fehr. Secure authentication from a weak key, without leaking information. Lecture Notes in Computer Science, 6632:246– 265, 2011. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer.com/content/
pdf/10.1007/978-3-642-20465-4_15.

Ahmadi:2011:SKC

[17] Hadi Ahmadi and Reihaneh Safavi-Naini.

Secret keys from channel noise. Lecture Notes in Computer Science, 6632:266– 283, 2011. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer.com/content/
pdf/10.1007/978-3-642-20465-4_16.

Obana:2011:AOC

[18] Satoshi Obana. Almost optimum t-cheater identifiable secret sharing schemes. Lecture Notes in Computer Science, 6632:284–302, 2011. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/content/pdf/10.1007/
978-3-642-20465-4_17.

Leander:2011:LHS

[19] Gregor Leander. On linear hulls, statistical saturation attacks, PRESENT and a cryptanalysis of PUFFIN. Lecture Notes in Computer Science, 6632:303– 322, 2011. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

                                        http://link.springer.com/content/
                                        pdf/10.1007/978-3-642-20465-4_18.

Dodis:2011:DEM

[20] Yevgeniy Dodis and John Steinberger. Domain extension for MACs beyond the birthday barrier. Lecture Notes in Computer Science, 6632:323–342, 2011. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

                                        /link.springer.com/content/pdf/10.
                                        1007/978-3-642-20465-4_19.

Sepehrdad:2011:SAR

[21] Pouyan Sepehrdad, Serge Vaudenay, and Martin Vuagnoux. Statistical attack on RC4. Lecture Notes in Computer Science, 6632:343–363, 2011. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer.
                                        com/content/pdf/10.1007/978-3-642-
                                        20465-4_20.

Becker:2011:IGA

[22] Anja Becker, Jean-Sébastien Coron, and Antoine Joux. Improved generic algorithms for hard knapsacks. Lecture Notes in Computer Science, 6632:364– 385, 2011. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

                                        http://link.springer.com/content/
                                        pdf/10.1007/978-3-642-20465-4_21.

shelat:2011:TOS

[23] Abhi shelat and Chih hao Shen. Twooutput secure computation with malicious adversaries. Lecture Notes in Computer Science, 6632:386–405, 2011. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

                                        /link.springer.com/content/pdf/10.
                                        1007/978-3-642-20465-4_22.

Ishai:2011:ENI

[24] Yuval Ishai, Eyal Kushilevitz, Rafail Ostrovsky, Manoj Prabhakaran, and Amit Sahai. Efficient non-interactive secure computation. Lecture Notes in Computer Science, 6632:406–425, 2011. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/content/pdf/10.
1007/978-3-642-20465-4_23.

Asharov:2011:TGT

[25] Gilad Asharov, Ran Canetti, and Carmit Hazay. Towards a game theoretic view of secure computation. Lecture Notes in Computer Science, 6632:426– 445, 2011. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer.com/content/
pdf/10.1007/978-3-642-20465-4_24.

Lindell:2011:HEU

[26] Yehuda Lindell. Highly-efficient universallycomposable commitments based on the DDH assumption. Lecture Notes in Computer Science, 6632:446–466, 2011. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/content/pdf/10.
1007/978-3-642-20465-4_25.

Unruh:2011:CCB

[27] Dominique Unruh. Concurrent composition in the bounded quantum storage model. Lecture Notes in Computer Science, 6632:467–486, 2011. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/content/pdf/10.1007/
978-3-642-20465-4_26.

Ristenpart:2011:CCL

[28] Thomas Ristenpart, Hovav Shacham, and Thomas Shrimpton. Careful with composition: Limitations of the indifferentiability framework. Lecture Notes in Computer Science, 6632:487–506, 2011. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/content/pdf/10.
1007/978-3-642-20465-4_27.

Malkin:2011:ECS

[29] Tal Malkin, Isamu Teranishi, and Moti Yung. Efficient circuit-size independent public key encryption with KDM security. Lecture Notes in Computer Science, 6632:507–526, 2011. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.
com/content/pdf/10.1007/978-3-642-
20465-4_28.

Applebaum:2011:KDM

[30] Benny Applebaum. Key-dependent message security: Generic amplification and completeness. Lecture Notes in Computer Science, 6632:527–546, 2011. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/content/pdf/10.
1007/978-3-642-20465-4_29.

Lewko:2011:UHA

[31] Allison Lewko and Brent Waters. Unbounded HIBE and attribute-based encryption. Lecture Notes in Computer Science, 6632:547–567, 2011. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/content/pdf/10.1007/
978-3-642-20465-4_30.

Lewko:2011:DAB

[32] Allison Lewko and Brent Waters. Decentralizing attribute-based encryption. Lecture Notes in Computer Science, 6632:568–588, 2011. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.
com/content/pdf/10.1007/978-3-642-
20465-4_31.

Wee:2011:TRC

[33] Hoeteck Wee. Threshold and revocation cryptosystems via extractable hash proofs. Lecture Notes in Computer Science, 6632:589–609, 2011. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/content/pdf/10.1007/
978-3-642-20465-4_32.

Durmuth:2011:DEN

[34] Markus Dürmuth and David Mandell Freeman. Deniable encryption with negligible detection probability: An interactive construction. Lecture Notes in Computer Science, 6632:610–626, 2011. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/content/pdf/10.
1007/978-3-642-20465-4_33.

Anonymous:2011:BM

[35] Anonymous. Back matter. Lecture Notes in Computer Science, 6632:??, 2011. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/content/pdf/bbm:
978-3-642-20465-4/1.

Kopetz:2011:WLB

[36] Hermann Kopetz. What I learned from Brian. Lecture Notes in Computer Science, 6875:1–6, 2011. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-
3-642-24541-1_1/;   http://link.
springer.com/content/pdf/bfm:978-
3-642-24541-1/1.pdf.

Lloyd:2011:BRB

[37] John L. Lloyd and Tom Anderson. Brian Randell: a biographical note. Lecture Notes in Computer Science, 6875:7–14, 2011. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer.com/chapter/
10.1007/978-3-642-24541-1_2/.

Aviienis:2011:BRA

[38] Algirdas Avi ienis. On building a referee’s avatar. Lecture Notes in Computer Science, 6875:15–22, 2011. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/
978-3-642-24541-1_3/.

Campbell-Kelly:2011:TPI

[39] Martin Campbell-Kelly. From theory to practice: The invention of programming, 1947–51. Lecture Notes in Computer Science, 6875:23–37, 2011. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/
978-3-642-24541-1_4/.

Lomet:2011:TLA

[40] David Lomet. Transactions: From local atomicity to atomicity in the cloud. Lecture Notes in Computer Science, 6875:38– 52, 2011. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer.com/chapter/
10.1007/978-3-642-24541-1_5/.

Rushby:2011:DM

[41] John Rushby. From DSS to MILS. Lecture Notes in Computer Science, 6875:53– 57, 2011. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer.com/chapter/
10.1007/978-3-642-24541-1_6/.

Swade:2011:PEC

[42] Doron Swade. Pre-electronic computing. Lecture Notes in Computer Science, 6875:58–83, 2011. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-
3-642-24541-1_7/.

Wichmann:2011:WW

[43] Brian Wichmann. Whetstone wanderings. Lecture Notes in Computer Science, 6875:84–92, 2011. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-
3-642-24541-1_8/.

Bacon:2011:URT

[44] Jean Bacon, Andrei Iu. Bejan, Alastair R.

Beresford, David Evans, and Richard J. Gibbens. Using real-time road traffic data to evaluate congestion. Lecture Notes in Computer Science, 6875:93–117, 2011. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer.com/chapter/
10.1007/978-3-642-24541-1_9/.

Bantre:2011:FTA

[45] Jean-Pierre Banâtre, Christine Morin, and Thierry Priol. Fault tolerant autonomic computing systems in a chemical setting. Lecture Notes in Computer Science, 6875:118–129, 2011. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/
978-3-642-24541-1_10/.

Bell:2011:CEY

[46] Gordon Bell. Out of a closet: The early years of The Computer Museum. Lecture Notes in Computer Science, 6875:130– 146, 2011. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer.com/chapter/
10.1007/978-3-642-24541-1_11/.

Burns:2011:TFM

[47] Alan Burns and Sanjoy Baruah. Timing faults and mixed criticality systems. Lecture Notes in Computer Science, 6875:147–166, 2011. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-
3-642-24541-1_12/.

Ceruzzi:2011:PBR

[48] Paul E. Ceruzzi. Professor Brian Randell and the history of computing. Lecture Notes in Computer Science, 6875:167– 173, 2011. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer.com/chapter/
10.1007/978-3-642-24541-1_13/.

Coffman:2011:CSF

[49] Ed Coffman. Computer storage fragmentation: Pioneering work of Brian Randell. Lecture Notes in Computer Science, 6875:174–184, 2011. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-
3-642-24541-1_14/.

Conway:2011:IAR

[50] Lynn Conway. IBM-ACS: Reminiscences and lessons learned from a 1960’s supercomputer project. Lecture Notes in Computer Science, 6875:185–224, 2011. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/
978-3-642-24541-1_15/.

Courtois:2011:BEM

[51] Pierre-Jacques Courtois. The Belgian Electronic Mathematical Machine (1951– 1962): an account. Lecture Notes in Computer Science, 6875:225–237, 2011. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/
978-3-642-24541-1_16/.

Dacier:2011:RDF

[52] Marc Dacier. On the resilience of the dependability framework to the intrusion of new security threats. Lecture Notes in Computer Science, 6875:238– 250, 2011. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer.com/chapter/
10.1007/978-3-642-24541-1_17/.

Denning:2011:VFT

[53] Peter J. Denning. Virtual fault tolerance. Lecture Notes in Computer Science, 6875:251–260, 2011. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-
3-642-24541-1_18/.

Hoare:2011:RB

[54] Tony Hoare. Recovery blocks. Lecture Notes in Computer Science, 6875:261– 266, 2011. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

                                        http://link.springer.com/chapter/
                                        10.1007/978-3-642-24541-1_19/.

Horning:2011:DWP

[55] J. J. Horning. The development and writing of “process structuring”. Lecture Notes in Computer Science, 6875:267– 272, 2011. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

                                        http://link.springer.com/chapter/
                                        10.1007/978-3-642-24541-1_20/.

Kiayias:2011:TCC

[56] Aggelos Kiayias, editor. Topics in cryptology — CT-RSA 2011: the cryptographers’ track at the RSA conference 2011, San Francisco, CA, USA, February 14– 18, 2011. proceedings, volume 6558 of Lecture notes in computer science. Springer-Verlag Inc., New York, NY, USA, 2011. ISBN 3-642-19073-1. LCCN ????

vanDam:2011:TQC

[57] Wim van Dam, Vivian M. Kendon, and Simone Severini, editors. Theory of quantum computation, communication, and cryptography: 5th conference, TQC 2010, Leeds, UK, April 13–15, 2010, revised selected papers, volume 6519 of Lecture notes in computer science. Springer-Verlag Inc., New York, NY, USA, 2011. ISBN 3-642-18072-8 (softcover). LCCN ????